# The AI Productivity Index: APEX-v1-extended

### v4

@@ Abstract
We present an extended version of the AI Productivity Index (APEX-v1-extended), a benchmark for assessing whether frontier models are capable of performing economically valuable tasks in four jobs: investment banking associate, management consultant, big law associate, and primary care physician (MD). This technical report details the extensions to APEX-v1, including an increase in the held-out evaluation set from $n=50$ to $n=100$ cases per job ($n=400$ total) and updates to the grading methodology. We present a new leaderboard, where GPT5 (Thinking = High) remains the top performing model with a score of $67.0\%$. APEX-v1-extended shows that frontier models still have substantial limitations when performing typical professional tasks. To support further research, we are open sourcing $n=25$ non-benchmark example cases per role ($n=100$ total) along with our evaluation harness.
@@ 1 Introduction
APEX-v1-extended contains realistic and varied prompts that reflect the everyday work of professionals. They are split across the same four knowledge jobs as the original v1 benchmark: investment banking associate, management consultant, big law associate, and primary care physician (MD). We have increased the size of the hidden heldout set from $n=50$ to $n=100$ per job, increasing the overall size of the benchmark from $n=200$ to $n=400$. The new cases have the same design, comprising a prompt, source documents that the model has to reason over, and a prompt-specific grading rubric. See an example in Figure  3. The $n=400$ cases will remain a closed heldout dataset for rigorous evaluation of frontier models. We are also open sourcing a dataset of $n=25$ cases per job ($n=100$ total) with a CC-BY license. We refer to this as apex-v1-devset. [^1] APEX-v1-extended on Hugging Face. Our eval harness is available on Github. [^2] Mercor on Github.
@@ 2.2 Quality control
Experts created each prompt independently, based on their day-to-day work. The mean time for an expert to execute the prompts in a real-world professional setting is estimated at $2.7$ hours, with a range of $0.5$ to $20$ hours. Experts submitted their prompts for review, where they were approved or rejected by reviewers. For rejections, reviewers could either remove the prompt from production or request specific changes. After the prompt was approved, the contributor created the grading rubric, which was also reviewed by a reviewer. Multiple rounds of review help to ensure that cases meet the requirements of the project, are varied in terms of topic, style and model behavior, and are sufficiently challenging. We use in-house LM-powered reviewing tools to give experts immediate feedback on their work. Experts are always responsible for the quality of their submissions. Finally, to ensure that models give a valid single-turn response, we append a short instruction to each prompt, as given in Appendix B.
@@ 2.4 Rubric creation
To scalably grade each model response, experts created a rubric of quality criteria (Saad-Falcon et al., 2024; Arora et al., 2025; Starace et al., 2025). Rubrics decompose the hard-to-measure concept of response “quality” into testable discrete components. Each criterion is an objective, specific, and self-contained statement about the response, phrased as a descriptive claim. They are analogous to unit tests for code, and can be assessed as Pass or Fail by either a judge LM or a human reviewer. The mean number of criteria per case is $14.81$ in APEX-v1-extended, compared with $11.40$ in APEX-v1-devset. This is a reduction from $29.09$ criteria in APEX-v1.0, which is due to (1) new cases being added with fewer criteria and (2) excess criteria being removed from the original cases.
@@ 3.2 Grading responses with an LM judge
LM judges are widely used to scalably assess the quality of model responses, supporting high-fidelity evals that can be re-run with no marginal human work (Baumann et al., 2025; Gu et al., 2025; Zhu et al., 2025). LM judging is well-suited for grading rubrics as each criterion is a short self-contained statement. In principle, these statements are easier to judge than overall quality or abstract concepts like “usefulness” or “helpfulness”. We used a single judge in APEX-v1-extended (Gemini 2.5 Flash) rather than the panel of judges for APEX-v1.0. This reduces the eval time, is more transparent, and maintains high quality. The prompt used by the LM judge remains the same, as given in Appendix A.
@@ 4.1 Leaderboard results
GPT 5 (Thinking = High) has the highest mean score at $67.0\%$, followed by Gemini 3 Pro (Thinking = High) at $64.3\%$ and Grok 4 at $63.5\%$. Models’ mean scores on APEX-v1-extended are shown in Figure 1 with $95\%$ confidence intervals. The differences between models’ scores are statistically significant overall using a Friedman omnibus test ($p<0.000001$). Models’ scores are lowest on Investment banking, with the highest scoring model achieving $63.0\%$, followed by Management consulting (top score = $64.0\%$), Medicine (top score = $65.5\%$), and, with substantially higher scores, Law (top score = $77.9\%$). Despite jobs’ differing difficulty, and the different knowledge and reasoning skills they require, models’ rank positions are fairly consistent. The three OpenAI models are particularly strong in law, taking the three top spots.
@@ 4.2 Performance given task difficulty
4.3 Comparison with APEX-v1-devset
We evaluate the same 10 models against the $n=100$ cases in APEX-v1-devset (available open source) using the exact same grading methodology (i.e., 8 runs and Gemini 2.5 Flash as a judge). Results are shown in Table 4. Overall, the open source data is similar but slightly more challenging. GPT 5 (Thinking = High) is the best scoring model on both sets, achieving $65.70\%$ on the APEX-v1-devset versus $67.0\%$ on APEX-v1-extended (difference = $1.3\%$). Differences in model ranks are minor (two sets of two models are swapped one position) and only one model performs better on the dev set than the benchmark (Gemini 3 Pro (Thinking = High) scoring $1.2\%$ higher). Most models find APEX-v1-devset harder, with GPT 5.1 (Thinking = High), Gemini 2.5 Flash (Thinking = On) and Sonnet 4.5 (Thinking = On) scoring lower by at least 10 percentage points. Finally, over the 8 runs, the mean standard deviation for APEX-v1-devset is slightly higher, increasing from $9.05$ to $10.76$.
@@ 5 Acknowledgments
We are grateful to all the expert annotators who contributed to APEX. We thank all our advisors and partners who gave feedback in shaping the design and release of APEX-v1-extended.

### v5

@@ Abstract
We present an extended version of the AI Productivity Index (APEX-v1-extended), a benchmark for assessing whether frontier models are capable of performing economically valuable tasks in four jobs: investment banking associate, management consultant, big law associate, and primary care physician (MD). This technical report details the extensions to APEX-v1, including an increase in the held-out evaluation set from $n=50$ to $n=100$ cases per job ($n=400$ total) and updates to the grading methodology. We present a new leaderboard, where GPT5 (Thinking = High) remains the top performing model with a score of $67.0\%$. APEX-v1-extended shows that frontier models still have substantial limitations when performing typical professional tasks. To support further research, we are open sourcing $n=25$ non-benchmark example cases per job ($n=100$ total) along with our evaluation harness.
@@ 1 Introduction
APEX-v1-extended contains realistic and varied prompts that reflect the everyday work of professionals. They are split across the same four knowledge jobs as the original v1 benchmark: investment banking associate, management consultant, big law associate, and primary care physician (MD). We have increased the size of the hidden heldout set from $n=50$ to $n=100$ per job, increasing the overall size of the benchmark from $n=200$ to $n=400$. The new cases have the same design, comprising a prompt, source documents that the model has to reason over, and a prompt-specific grading rubric. See an example in Figure  3. The $n=400$ cases will remain a closed heldout dataset for rigorous evaluation of frontier models. We are also open sourcing a dataset of $n=25$ cases per job ($n=100$ total) with a $C$-BY license. We refer to this as apex-v1-devset. [^1] APEX-v1-extended on Hugging Face. Our eval harness is available on Github. [^2] Mercor on Github.
@@ 2.2 Quality control
Experts created each prompt independently, based on their day-to-day work. The mean time for an expert to execute the prompts in a real-world professional setting is estimated at $2.7$ hours, with a range of $0.5$ to 2020 hours. Experts submitted their prompts for review, where they were approved or rejected by reviewers. For rejections, reviewers could either remove the prompt from production or request specific changes. After the prompt was approved, the contributor created the grading rubric, which was also reviewed by a reviewer. Multiple rounds of review help to ensure that cases meet the requirements of the project, are varied in terms of topic, style and model behavior, and are sufficiently challenging. We use in-house LM-powered reviewing tools to give experts immediate feedback on their work. Experts are always responsible for the quality of their submissions. Finally, to ensure that models give a valid single-turn response, we append a short instruction to each prompt, as given in Appendix B.
@@ 2.4 Rubric creation
To scalably grade each model response, experts created a rubric of quality criteria (Saad-Falcon et al., 2024; Arora et al., 2025; Starace et al., 2025). Rubrics decompose the hard-to-measure concept of response “quality” into testable discrete components. Each criterion is an objective, specific, and self-contained statement about the response, phrased as a descriptive claim. They are analogous to unit tests for code, and can be assessed as Pass or Fail by either a judge LM or a human reviewer. The mean number of criteria per case is $14.81$ in APEX-v1-extended, compared with $11.40$ in APEX-v1-devset. This is a reduction from $29.09$ criteria in APEX-v1.0, due to (1) new cases being added with fewer criteria and (2) excess criteria being removed from the original cases.
@@ 3.2 Grading responses with an LM judge
LM judges are widely used to scalably assess the quality of model responses, supporting high-fidelity evals that can be re-run with no marginal human work (Baumann et al., 2025; Gu et al., 2025; Zhu et al., 2025). LM judging is well-suited for grading rubrics as each criterion is a short self-contained statement. In principle, these statements are easier to judge than overall quality or abstract concepts like “usefulness” or “helpfulness”. We used a single judge in APEX-v1-extended (Gemini 2.5 Flash) rather than the panel of judges for APEX-v1.0. This reduces the eval time, is more transparent, and maintains high performance. The prompt used by the LM judge remains the same, as given in Appendix A.
@@ 4.1 Leaderboard results
GPT 5 (Thinking = High) has the highest mean score at $67.0\%$, followed by Gemini 3 Pro (Thinking = High) at $64.3\%$ and Grok 4 at $63.5\%$. Models’ mean scores on APEX-v1-extended are shown in Figure 1 with $95\%$ confidence intervals. The differences between models’ scores are statistically significant overall using a Friedman omnibus test ($p<0.000001$). Models’ scores vary across jobs. They are lowest for Investment banking, with the highest scoring model achieving $63.0\%$, followed by Management consulting (top score = $64.0\%$), Medicine (top score = $65.5\%$), and, with substantially higher scores, Law (top score = $77.9\%$). Despite jobs’ differing difficulty, and the different knowledge and reasoning skills they require, models’ rank positions are fairly consistent.
@@ 4.2 Performance given task difficulty
4.3 Comparison of APEX-v1-extended and APEX-v1-devset
We evaluate the same 10 models against the $n=100$ cases in APEX-v1-devset using the exact same grading methodology (i.e., 8 runs and Gemini 2.5 Flash as a judge). Results are shown in Table 4. GPT 5 (Thinking = High) is the best scoring model on both sets, achieving $65.70\%$ on the APEX-v1-devset versus $67.0\%$ on APEX-v1-extended. Differences in model ranks are minor, with two sets of two models swapped one position, and only one model performs better on APEX-v1-devset than APEX-v1-extended (Gemini 3 Pro (Thinking = High) scoring $1.2\%$ higher). GPT 5.1 (Thinking = High), Gemini 2.5 Flash (Thinking = On) and Sonnet 4.5 (Thinking = On) all score lower on APEX-v1-devset by more than 10 percentage points.
@@ 5 Acknowledgments
We are grateful to the expert annotators who contributed to APEX. We thank all our advisors and partners who gave feedback.